### v1

# All-optical mapping of ultrafast carrier dynamics in a Dirac semimetal

Abstract
High-harmonic generation (HHG), the hallmark effect of attosecond science, is a nonperturbative nonlinear process leading to the emission of high harmonic light from gases and solids. In gases, extreme driving laser pulse intensities can deplete the ground state, suppressing harmonic emission during the trailing edge of the pulse. Here, we report pronounced ultrafast depletion dynamics during HHG in a gapless Dirac semimetal—highly oriented pyrolytic graphite (HOPG). Remarkably, HOPG supports nonperturbative harmonic generation at laser intensities as low as W cm-2, facilitated by its vanishing bandgap. Using two-color spectroscopy, we reveal the excitation dynamics of Dirac electron-hole pairs as it affects the emission of harmonics during the presence of the driving laser pulse. Notably, valence band depletion near the Dirac points leads to a marked suppression of interband harmonics and induces measurable temporal shifts. These observations are supported by simulations based on semiconductor Bloch equations. Our findings uncover a new regime of strong-field light–matter interaction in gapless solids and establish HHG as a sensitive, all-optical probe of ultrafast carrier dynamics, offering new opportunities for ultrafast optoelectronics in Dirac materials.
Main
Parametric nonlinear optical processes, such as second-harmonic generation or four-wave mixing, transiently modify the quantum state of a medium during the laser interaction, but leave it unchanged afterward 1. This framework extends even to high-harmonic generation (HHG), an extreme nonlinear strong-field-driven process observed mainly in gases 2, 3, 4 and condensed-matter systems 5, 6, 7, 8, 9, which cannot be described by perturbative nonlinear optics. Although HHG itself is parametric, competition with other strong-field processes can affect HHG, most prominently photoionization 4. In gases, sufficiently intense and long laser pulses can substantially deplete the ground-state electron population, suppressing harmonic emission, causing spectral shifts and reducing harmonic cutoff energies 10, 11. Additionally, the resulting high densities of photoelectrons adversely affect harmonic phase matching 12, 13.
In solids, analogous depletion phenomena might be expected at considerably lower laser intensities due to their smaller bandgaps relative to atomic ionization energies. However, experimental observation of pronounced depletion dynamics during solid-state HHG remains elusive. Achieving a significant suppression of harmonic emission typically requires excitation fractions on the order of , while in experiment, typical excitation levels in semiconductors are usually limited to a few percent to avoid material damage 14. Pre-excitation of carriers with ultraviolet pulses has been explored as an alternative route to reach the depletion regime and suppress the HHG yield in semiconductors 14, 15, 16. However, such approaches obscure the intrinsic depletion dynamics driven solely by the fundamental laser pulse which drives the HHG process. Here, the subject of depletion and HHG yield suppression is an intricate one because condensed matter is far more complex than the gas phase due to its strong multi-electron nature, lower symmetry, and Bloch state delocalization. Moreover, the microscopic origins of harmonic suppression in solids remain actively debated. Suppression could arise predominantly from Pauli blocking, where transitions from valence to conduction bands become unavailable due to occupied states, or from excitation-induced dephasing, where increased carrier densities accelerate electron scattering and degrade interband coherence 17.
Highly oriented pyrolytic graphite (HOPG), a structurally robust carbon-based bulk material composed of stacked graphene layers 18, 19, provides a promising platform for investigating such intrinsic depletion dynamics under strong-field driving. HOPG retains the key electronic properties of monolayer graphene, including its gapless Dirac dispersion, high carrier mobility, and strong anisotropic optical responses 20, 21, 22. It also offers additional experimental advantages, such as structural robustness and scalability, that overcome the intrinsic limitation of graphene’s low damage threshold. Furthermore, HOPG allows systematic exploration of the effects of interlayer coupling and thickness variations on nonlinear optical and electronic dynamics 23. Despite these compelling features, HHG in HOPG remains experimentally unexplored to date.
In this combined experimental and theoretical study, we introduce an all-optical two-color spectroscopy technique to track population dynamics in the Dirac cone, revealing behavior beyond that observed in conventional gapped solids. Specifically, we observe a time shift in harmonic emission relative to the driving field peak, a clear signature of strong-field-induced depletion in a gapless Dirac semimetal. While prior studies have investigated photon-induced saturation in graphene for applications, such as saturable absorbers 24, our work explores a fundamentally distinct regime, the nonperturbative strong-field limit, where carrier depletion evolves within a few optical cycles and directly influences HHG. This bridges the gap between weak-field saturation effects and strong-field dynamics in Dirac materials, uncovering a new class of ultrafast phenomena governed by light-driven population reshaping. More broadly, the ability to resolve femtosecond-scale carrier depletion through HHG suppression suggests that this approach could also serve as a sensitive, all-optical probe of light-induced phase transitions in solids 25, 26, 27.
Results
Nonperturbative harmonic generation in HOPG
We generate high harmonics from an HOPG sample in reflection geometry 28 using an infrared femtosecond pulse (frequency ) combined with a weak, synchronized second harmonic (SH) field (), with a polarization parallel to the fundamental pulse (see Fig. 1a). The fundamental pump pulse has a central wavelength of 1980 nm and a duration of 60 fs (about 10 optical cycles). The SH field is generated by focusing the fundamental field onto a -barium borate (BBO) crystal. Both beams are split and then recombined in an interferometric setup, achieving both spatial and temporal overlap with accurate control of the two-color (–) time delay (see Methods). Throughout this work, we set the SH pulse intensity at 3% of the fundamental pump intensity in both the experimental measurements and theory calculations.
HOPG exhibits an AB-stacked multilayer graphene structure, resulting in a stable crystalline form whose band dispersion near the Dirac point closely resembles that of monolayer graphene. For comparison, ab initio calculations of the electronic band structures of both graphene and HOPG are provided in Fig. S11 of the Supplementary Information. Given the moderate pump field strength used in our experiments – limited by the damage threshold of the material – the electron dynamics are effectively confined to the vicinity of the Dirac cone (Fig. 1b). In the nonperturbative regime relevant to our work, the emission of above-bandgap harmonics is primarily governed by interband polarization processes. This behavior is further corroborated by numerical simulations based on the semiconductor Bloch equations (SBEs), which reveal a clear predominance of interband contributions in the generated harmonic spectrum (see Fig. S5 in Supplementary Information). Interband polarization roughly corresponds to the three-step picture of HHG. Here, interband tunneling produces an electron-hole pair that is driven by the laser field in momentum space until electron and hole recombine and a harmonic photon is emitted 6.
In our experiment, non-perturbative high harmonics are generated from HOPG with spectral features observed up to the fourth order (2.5 eV photon energy) near the reported harmonic cutoff in monolayer graphene around 2.4 eV 29, 30. Figure 1c displays the measured harmonic spectrum obtained using the 1980 nm driving pulse with a peak intensity of 72 GW cm-2 and the admixed SH pulse. With respect to the third harmonic, the fourth harmonic exhibits a relative intensity of , while the second harmonic is omitted from the spectrum, as it is masked by the much more intense incident SH field.
To confirm the nonperturbative nature of the harmonic emission, we measure the intensity scaling of the third harmonic (H3) driven by a single-color 1980 nm pulse (Fig. 1d). The third harmonic exhibits a power-law dependence of approximately , deviating from the expected perturbative scaling of and confirming the involvement of strong-field electron-hole dynamics. This nonlinear response arises from electron-hole acceleration and recombination processes rather than multiphoton transitions, underscoring that our measurements are firmly within the non-perturbative regime.
Emission time shift of harmonic generation
Two-color spectroscopy is a powerful all-optical approach capable of retrieving the temporal dynamics of the harmonic emission all the way down to the sub-cycle scale 31. Here, we measure the harmonic intensity as a function of the relative time delay between the two colors (–), as shown in Fig. 2. The fourth harmonic (H4) is generated in HOPG when the and pulses temporally overlap, enabled by symmetry breaking of the fundamental electric field through perturbation by the pulse (Fig. 2a). As a reference, we replace the HOPG sample with a conventional bulk gapped solid HHG sample, a ZnO crystal 5, 6, and measure the above-band gap sixth harmonic (H6) intensity as a function of . The measurements in ZnO are performed under the same conditions as in HOPG, except that the fundamental pump intensity is increased to 384 GW cm-2 to ensure the generation of above-band gap harmonics through the interband process. The band gap of the ZnO crystal is around 3.3 eV. Full delay-dependent harmonic spectra are shown in Fig. S2 of the Supplementary Information.
In each of plots shown in Fig. 2, oscillations of the harmonic signal with a frequency of 4 are visible on top of an overall Gaussian-like curve. In our work, we focus on the latter. Remarkably, we observe that the maximum signal of H4 in HOPG occurs approximately fs before the temporal overlap of the two pulse intensity peaks at . In contrast, the H6 signal in ZnO exhibits a symmetric delay dependence, with its maximum occurring precisely at zero delay. The time shift corresponding to the maximum harmonic signal is extracted by fitting the entire delay-dependent harmonic trace with a Gaussian function and identifying the peak position of the fit. To gain further insight into the underlying physics, we perform theoretical simulations based on both one-dimensional (1D) and two-dimensional (2D) semiconductor Bloch equations (SBEs, see Methods in the main text and Supplementary Information for details) using the same parameters as in the experiment for both HOPG and ZnO (Fig. 2c and d; see also full spectra in Fig. S6 of the Supplementary Information). Under a fixed interband decoherence time of fs, both the 1D and 2D models exhibit a negative delay for the harmonic maximum (note that the negative delay in 2D simulation is sensitive to the choice of T2). Since the 1D SBE model reproduces the main features observed in the 2D simulations, including the negative time shift of the harmonic maximum, we adopt the simpler 1D model for the remainder of our analysis. The simulation results show excellent agreement with the experimental data. We accurately reproduce the time shift of the H4 signal in HOPG, yielding a similar shift fs with respect to , and correctly predict the peak of the H6 signal in ZnO at zero delay. Furthermore, we simulate the delay dependence of H4, H6, and H8 from ZnO, and in all cases, the harmonic maxima occur near , confirming that the choice of H6 as a reference does not affect the evaluation of the delay.
To gain deeper insight into the pronounced time shift observed in HHG from HOPG, we analyze the time evolution of both the conduction band electron population and the interband polarization , as shown in Figs. 3a and b. Our calculations reveal that this shift originates from the strong interplay between intraband and interband currents in the gapless Dirac cone. A significant accumulation of intraband population leads to a suppression of the interband polarization. Notably, the conduction band population near the Dirac point rapidly approaches half-filling well before the peak of the driving field, indicating a pronounced saturation effect. Unlike conventional solid-state HHG processes where depletion is negligible 32, the time-dependent -space populations of conduction () and valence () bands play a crucial role in the interband transition term in the SBEs (see the Methods section, Eq. (1)). Once both and reach 0.5, the interband transition term vanishes, effectively blocking further interband excitation – a phenomenon known as Pauli blocking. This behavior is analogous to a two-level system where population inversion is forbidden and the excitation process saturates due to balanced electron occupation across bands.
The inability of the conduction band population to significantly exceed half-filling for most -points around the Dirac point indicates that saturation occurs early during the fundamental driving pulse. This saturation not only limits further excitation, but also suppresses the interband polarization, which is an essential ingredient for interband HHG. As a result, the dominant interband term is weakened, even near the peak of the driving pulse. Instead, the maximum interband harmonic intensity occurs prior to the onset of saturation.
To visualize this effect more clearly, we compute the total conduction band population (around the cone) by integrating over momentum space, , shown in Fig. 3c. It is evident that the excitation of 50% of the final total population (which remains after the end of the laser pulse) is achieved at a significantly earlier time than the peak of the driving field, underscoring the presence of strong saturation and depletion. In the absence of depletion, one would expect this time to coincide with the peak of the driving field due to its temporal symmetry. For comparison, we also plot the calculated total conduction band population dynamics in ZnO, which shows that the excitation of 50% of the final total population occurs right at the peak field, consistent with the larger band gap that prevents efficient depletion.
Furthermore, a time-frequency analysis using a Gabor transform (Fig. 3d) shows that the maximum intensity of the H3 in HOPG driven by a single-color 2 m pulse occurs approximately 25 fs before the field maximum (). A complementary Gabor transform of the harmonic emission driven by – two-color pulses at various time delays (see Fig. S8 in the Supplementary Information) demonstrates that the field primarily acts as a probe, sampling the harmonic emission originating from the single-color excitation. This is in contrast to wide-band gap semiconductors where the non-depletion approximation is valid and interband polarization is largely unaffected by intraband population. In our case, the absence of a band gap in HOPG leads to substantial carrier excitation around the Dirac cone and strongly modifies the interband contribution to HHG. Consequently, intraband population dynamics play a critical role in shaping the overall HHG process in gapless Dirac materials such as HOPG. One can also expect to see a strong saturation effect in wide-bandgap semiconductor HHG as long as the pump intensity is sufficiently large. However, the low damage threshold of these materials prevents a clear observation of this effect in gapped solids. Nonetheless, our simulations confirm that the harmonic emission time shifts significantly earlier even in ZnO when its band gap is artificially reduced in simulations to a nearly gapless regime, as shown in Fig. S9 of the Supplementary Information.
Since the emission time shift originates from dynamic population saturation, a stronger driving laser intensity is expected to induce earlier saturation during the driving process. To investigate this, we experimentally measure the time shift of H4 under various pump intensities, as shown in Fig. 4a. As the driving laser intensity increases, the peak of the H4 signal in the – delay scan systematically shifts to earlier times, further corroborating the presence of a saturation effect. The relationship between the H4 emission time shift and the driving laser intensity is plotted in Fig. 4b.
Our SBE-based simulations accurately reproduce the experimental trend, confirming that the negative time shift increases with increasing peak intensity. To gain further insight into the phenomenon, we introduce an analytical approximation (Eqs. 8–12) based on Fermi’s golden rule, assuming non-depleted excitation from the valence band to the conduction band while accounting for the influence of electron occupation in both bands on the excitation rate. This model also predicts a similar dependence of on the pump intensity, as indicated by the black dashed line in Fig. 4b.
Dirac electron population dynamics
Because the time shift of the maximum harmonic intensity induced by saturation is highly sensitive to the electron population, it can serve as an effective probe for tracking electron population dynamics. To investigate this, we analyze the electron dynamics associated with H4 generation through the interband process by integrating the conduction band population over the momentum-space region where the bandgap energy corresponds to the H4 photon energy (see Fig. 5a).
Under excitation by a 2 m pulse with a peak intensity of 70 GW cm-2, the electron population contributing to H4 exhibits strong temporal oscillations (see Fig. 5b). We define as the time at which half of the electrons in this momentum-space region have transitioned to the conduction band. As shown in Fig. 5c, the absolute value of increases with pump intensity, indicating an intensity-dependent excitation process. Furthermore, a strong correlation is observed between (in the single-color 2 m excitation case) and (emission time shift corresponding to the maximum H4 intensity in the two-color excitation case). The correspondence can be captured by a second-order polynomial fitting function, , with an value of 0.99, indicating very strong correlation between these two physical observables. This demonstrates an all-optical method for probing carrier population dynamics in Dirac materials and semi-metallic systems under strong-field excitation.
Conclusion
In conclusion, our joint experimental and theoretical study has established an HHG-based all-optical approach to observe and quantify ultrafast depletion effects and carrier excitation dynamics in solids, which also yields absolute timing information. With respect to conventional bulk materials, such as ZnO, we found that the Dirac semimetal HOPG shows strong valence band population saturation due to its gapless nature even under moderate laser intensities (), leading to half-filling in the vicinity of the Dirac point. We identified this effect as responsible for the observed fs time shifts and the suppression of the harmonic emission in the later part of the driving laser pulse through Pauli blocking.
Our study demonstrates that gapless materials are subject to depletion effects and Pauli blocking, leaving behind a sizable conduction band population and rendering harmonic generation processes nonparametric in nature. Petahertz optoelectronics, where optical manipulation of the material is performed on attosecond time scales, relies mainly on virtual charge carrier excitations and their reversibility 33, 34, 8. In order to realize petahertz optoelectronics with full reversibility in gapless materials, pulse duration and peak intensity need to be limited such that Pauli blocking is avoided. Even if this cannot be achieved, gapless materials can nevertheless serve as an ultrafast optoelectronic switch where the presence or the absence of a pre-excitation pulse affects the efficiency of nonlinear processes through Pauli blocking on the significantly longer time scale of spontaneous emission. The nonparametric nature of HHG in gapless materials even at moderate intensities may also be helpful for generating upconverted light with nonclassical photon statistics due to the depletion effect, as recently predicted for gas-phase HHG 35.
Methods
Experimental setup
In the experiment, the fundamental () pulse at 1980 nm is generated from a phase-stable optical parametric chirped pulse amplification (OPCPA) laser system operating at a repetition rate of 1 MHz (Class 5 Photonics White Dwarf 36). The pulse duration of the field measured by frequency-resolved optical gating (FROG) yields a full-width-at-half-maximum (FWHM) duration of approximately 60 fs ( optical cycles). The second harmonic () pulse was generated by focusing the fundamental beam into a -barium borate (BBO) crystal under optimum phase-matching conditions. By adjusting the power, we maintained the intensity ratio between the and pulses at . For HHG in HOPG, the pulse energy was set between 1–5 nJ, while for ZnO, a higher energy of approximately 30 nJ was used.
The two-color interferometer setup was implemented by first separating the and beams using a dichroic beam splitter. The relative delay between the two pulses was controlled via a piezo delay stage placed in the beam path. The two beams were then recombined by reflecting them off two mirrors back into the same dichroic beam splitter, with a slight lateral offset angle relative to the incoming paths, and subsequently directed toward the sample for HHG. A parabolic mirror with mm focuses the light onto the HOPG sample (TipsNano) at an angle of incidence of 14 degrees. The beam size at the focus was measured using a knife-edge method, yielding an effective radius of (see discussion and Fig. S1 in the Supplementary Information). We collect the harmonic light with a lens and send it to a Si-based spectrometer.
During the two-color delay scan, the piezo stage exhibited a minimal drift of less than 1 fs, which does not affect the accuracy of the delay-dependent measurements. Furthermore, to actively monitor and correct for any such drift, a portion of the combined – beam was directed into an additional BBO crystal to generate a reference SH signal. The resulting modulation arises from the interference between the newly generated SH signal from the beam and the original pulse. This delay-dependent modulation serves as a timing reference, allowing for post-correction of any small delay drifts in the piezo stage.
The thickness of the reference ZnO crystal (FarView, ) is 0.5 mm. Since the harmonics are measured in a reflection geometry on the crystal surface, nonlinear propagation effects associated with the crystal thickness do not influence our results.
Semiconductor Bloch equations calculations
To describe the interaction of the two-color field and HOPG in the Dirac cone, we use the following one-dimensional semiconductor Bloch equations37, 38 (SBEs) centered around the Dirac cone:
Here, represents a dimensionless polarization depending on time and momentum , induced by the time-dependent electric field . The indices and correspond to the valence band and conduction band, respectively. denotes the population of electrons (holes) in the conduction (valence) band. Here, we consider linear energy-momentum dispersion relation around Dirac cone with Fermi velocity describing a continuum ground state Dirac Hamiltonian. This approximation should be valid in our regime where the laser-driven electrons only move near the Dirac cone. The transition dipole moments (TDMs) are calculated as 39 (see Fig. S3 in the Supplementary Information). We set the phenomenological interband polarization decoherence time to one optical cycle of the 2 m driving laser field fs, and the relaxation time for intraband current to . This consideration is valid since the relaxation is generally long and ranges from few hundred femtosecond to picosecond in graphene and is less sensitive to the results 40, 41. Band structure and TDM for ZnO crystal along the –M direction are shown in Fig. S4 of the Supplementary Information (for details see the Supplementary Information). The calculations are performed with dense -grid of size 200, using a time step of 2.4 as and according to the experimental parameters. The yield of the harmonics for the two-color case are calculated with two-color delay steps of 0.33 fs.
The HHG spectrum is computed as 37
where
and
denote the interband and intraband contributions to the HHG yield, respectively. represents the group velocity of the electrons and holes in the corresponding bands.
Analytical form of population dynamics
For a theoretical analysis of saturation and electron population dynamics, we begin by considering the non-depleted excitation rate of electrons from the valence band to the conduction band. According to Fermi’s golden rule, this excitation rate at a given crystal momentum and time can be expressed as
The total excitation rate integrated over the Brillouin zone is then given by
Let denote the total conduction band electron population at time . The evolution of this population in discrete time steps can be described by
Transitioning to the continuous-time limit, we obtain the differential form
Integrating yields the electron population as a function of time,
The actual (depleted) instantaneous excitation rate is given by
For interband-driven high-harmonic generation (HHG), the emitted harmonic intensity is proportional to the instantaneous excitation rate, .
Data availability
The data that support the plots within this paper and other findings of this study are available from the authors upon request.
Code availability
The codes that support the findings of this study are available from the authors upon request.
Author contributions
Z. C. and C. G. contributed equally to this work.
M. K. and M. F. C. conceived and supervised the project, joined by O. N. in the later stage of the project. Z. C. built the experimental setup and performed the measurements. C. G. performed initial simulations. Z. C. carried out the theory calculations and C. G. and O. N. performed independent supplemental calculations. Z. C. and C. G. wrote the initial manuscript. All authors contributed to the interpretation of the results and to the preparation of the final manuscript.
Corresponding authors
Correspondence to Michael Krüger or Marcelo F. Ciappina.
Acknowledgements.
We thank M. Ivanov, M. Segev, Q. Yan and Y. Wu for insightful discussions and U. Leonhardt and L. M. Procopio for providing specialized equipment. Z. C., C. G., I. N., D. K., M. F. C and M. K. acknowledge the Guangdong Technion – Israel Institute of Technology (GTIIT) and Technion Seed Grant Program for enabling their collaboration and joint research. Z. C., I. N., D. K., M. F. C and M. K. thank the Helen Diller Quantum Center at the Technion for partial financial support. M. F. C. acknowledges support by the National Key Research and Development Program of China (Grant No. 2023YFA1407100), Guangdong Province Science and Technology Major Project (Future functional materials under extreme conditions - 2021B0301030005) and the Guangdong Natural Science Foundation (General Program project No. 2023A1515010871).Ethics declarations
The authors declare no competing interests.

### v2

@@
-All-optical mapping of ultrafast carrier dynamics in a Dirac semimetal
+Two-color harmonic spectroscopy of ultrafast Dirac electron dynamics
 Abstract
-High-harmonic generation (HHG), the hallmark effect of attosecond science, is a nonperturbative nonlinear process leading to the emission of high harmonic light from gases and solids.
+High-harmonic generation (HHG), the hallmark effect of attosecond science, is a nonperturbative nonlinear process leading to the emission of high-harmonic light from gases and solids.
 In gases, extreme driving laser pulse intensities can deplete the ground state, suppressing harmonic emission during the trailing edge of the pulse.
-Here, we report pronounced ultrafast depletion dynamics during HHG in a gapless Dirac semimetal—highly oriented pyrolytic graphite (HOPG).
-Remarkably, HOPG supports nonperturbative harmonic generation at laser intensities as low as W cm-2, facilitated by its vanishing bandgap.
+Here, we report a similar effect, pronounced ultrafast carrier saturation dynamics and harmonic emission suppression during nonperturbative harmonic generation (NPHG) in a gapless Dirac semimetal—highly oriented pyrolytic graphite (HOPG).
+Remarkably, HOPG supports NPHG at laser intensities as low as W cm-2, facilitated by its vanishing bandgap.
+Ultrafast carrier saturation strongly modulates the interplay between interband and intraband currents, a key characteristic of NPHG in Dirac materials.
 Using two-color spectroscopy, we reveal the excitation dynamics of Dirac electron-hole pairs as it affects the emission of harmonics during the presence of the driving laser pulse.
-Notably, valence band depletion near the Dirac points leads to a marked suppression of interband harmonics and induces measurable temporal shifts.
+The excitation of out-of-equilibrium hot carriers and the concomitant saturation near the Dirac points leads to a marked suppression of interband harmonics and induces measurable temporal shifts.
 These observations are supported by simulations based on semiconductor Bloch equations.
-Our findings uncover a new regime of strong-field light–matter interaction in gapless solids and establish HHG as a sensitive, all-optical probe of ultrafast carrier dynamics, offering new opportunities for ultrafast optoelectronics in Dirac materials.
+Our finding reveal that field-driven carrier saturation plays a critical role in gapless solid NPHG.
+We demonstrate the potential of NPHG and HHG as a sensitive, all-optical probe of ultrafast carrier dynamics, offering novel opportunities for ultrafast optoelectronics in Dirac materials.
 Main
 Parametric nonlinear optical processes, such as second-harmonic generation or four-wave mixing, transiently modify the quantum state of a medium during the laser interaction, but leave it unchanged afterward 1.
-This framework extends even to high-harmonic generation (HHG), an extreme nonlinear strong-field-driven process observed mainly in gases 2, 3, 4 and condensed-matter systems 5, 6, 7, 8, 9, which cannot be described by perturbative nonlinear optics.
+This framework extends even to high-harmonic generation (HHG), an extreme nonlinear field-driven process observed mainly in gases 2, 3, 4 and condensed-matter systems 5, 6, 7, 8, 9, 10, which cannot be described by perturbative nonlinear optics.
 Although HHG itself is parametric, competition with other strong-field processes can affect HHG, most prominently photoionization 4.
-In gases, sufficiently intense and long laser pulses can substantially deplete the ground-state electron population, suppressing harmonic emission, causing spectral shifts and reducing harmonic cutoff energies 10, 11.
-Additionally, the resulting high densities of photoelectrons adversely affect harmonic phase matching 12, 13.
-In solids, analogous depletion phenomena might be expected at considerably lower laser intensities due to their smaller bandgaps relative to atomic ionization energies.
-However, experimental observation of pronounced depletion dynamics during solid-state HHG remains elusive.
-Achieving a significant suppression of harmonic emission typically requires excitation fractions on the order of , while in experiment, typical excitation levels in semiconductors are usually limited to a few percent to avoid material damage 14.
-Pre-excitation of carriers with ultraviolet pulses has been explored as an alternative route to reach the depletion regime and suppress the HHG yield in semiconductors 14, 15, 16.
-However, such approaches obscure the intrinsic depletion dynamics driven solely by the fundamental laser pulse which drives the HHG process.
-Here, the subject of depletion and HHG yield suppression is an intricate one because condensed matter is far more complex than the gas phase due to its strong multi-electron nature, lower symmetry, and Bloch state delocalization.
+In gases, sufficiently intense and long laser pulses can substantially deplete the ground-state electron population, suppressing harmonic emission, causing spectral shifts and reducing harmonic cutoff energies 11, 12.
+Additionally, the resulting high densities of photoelectrons adversely affect harmonic phase matching 13, 14.
+In solids, analogous phenomena causing the suppression of the emission of harmonics might be expected at considerably lower laser intensities due to their smaller bandgaps relative to atomic ionization energies.
+However, experimental observation of pronounced carrier excitation dynamics during solid-state HHG and, more generally, nonperturbative harmonic generation (NPHG) remains elusive.
+Achieving a significant suppression of harmonic emission typically requires excitation fractions on the order of , while in experiment, typical excitation levels in semiconductors are usually limited to a few percent to avoid material damage 15.
+Pre-excitation of carriers with ultraviolet pulses has been explored as an alternative route to reach a carrier saturation regime and suppress the harmonic yield in semiconductors 15, 16, 17.
+However, such approaches obscure the intrinsic saturation dynamics driven solely by the fundamental laser pulse which drives the harmonic generation process.
+Here, the subject of saturation and harmonic yield suppression is an intricate one because condensed matter is far more complex than the gas phase due to many-body scattering and decoherence, lower symmetry, and Bloch state delocalization.
 Moreover, the microscopic origins of harmonic suppression in solids remain actively debated.
-Suppression could arise predominantly from Pauli blocking, where transitions from valence to conduction bands become unavailable due to occupied states, or from excitation-induced dephasing, where increased carrier densities accelerate electron scattering and degrade interband coherence 17.
-Highly oriented pyrolytic graphite (HOPG), a structurally robust carbon-based bulk material composed of stacked graphene layers 18, 19, provides a promising platform for investigating such intrinsic depletion dynamics under strong-field driving.
-HOPG retains the key electronic properties of monolayer graphene, including its gapless Dirac dispersion, high carrier mobility, and strong anisotropic optical responses 20, 21, 22.
+Such suppression could arise predominantly from state blocking, where transitions from valence to conduction bands become unavailable due to occupied states, or from excitation-induced dephasing, where increased carrier densities accelerate electron scattering and degrade interband coherence 18.
+Highly oriented pyrolytic graphite (HOPG), a structurally robust carbon-based bulk material composed of stacked graphene layers 19, 20, provides a promising platform for investigating such intrinsic saturation dynamics under laser-field driving.
+HOPG retains the key electronic properties of monolayer graphene, including its gapless Dirac dispersion, high carrier mobility, and strong anisotropic optical responses 21, 22, 23.
 It also offers additional experimental advantages, such as structural robustness and scalability, that overcome the intrinsic limitation of graphene’s low damage threshold.
-Furthermore, HOPG allows systematic exploration of the effects of interlayer coupling and thickness variations on nonlinear optical and electronic dynamics 23.
-Despite these compelling features, HHG in HOPG remains experimentally unexplored to date.
-In this combined experimental and theoretical study, we introduce an all-optical two-color spectroscopy technique to track population dynamics in the Dirac cone, revealing behavior beyond that observed in conventional gapped solids.
-Specifically, we observe a time shift in harmonic emission relative to the driving field peak, a clear signature of strong-field-induced depletion in a gapless Dirac semimetal.
-While prior studies have investigated photon-induced saturation in graphene for applications, such as saturable absorbers 24, our work explores a fundamentally distinct regime, the nonperturbative strong-field limit, where carrier depletion evolves within a few optical cycles and directly influences HHG.
-This bridges the gap between weak-field saturation effects and strong-field dynamics in Dirac materials, uncovering a new class of ultrafast phenomena governed by light-driven population reshaping.
-More broadly, the ability to resolve femtosecond-scale carrier depletion through HHG suppression suggests that this approach could also serve as a sensitive, all-optical probe of light-induced phase transitions in solids 25, 26, 27.
+Furthermore, HOPG allows systematic exploration of the effects of interlayer coupling and thickness variations on nonlinear optical and electronic dynamics 24.
+Despite these compelling features, HHG and NPHG of low-order harmonics in HOPG remain experimentally unexplored to date.
+In this combined experimental and theoretical study, we introduce an all-optical two-color spectroscopy technique to study ultrafast population dynamics in the Dirac cone, revealing behavior beyond that observed in conventional gapped solids.
+Specifically, we observe a time shift in harmonic emission relative to the driving field peak, a clear signature of laser-field-induced carrier saturation in a gapless Dirac semimetal.
+Whereas earlier studies of graphite and graphene have probed photon-induced saturation, Pauli blocking, and relaxation dynamics using transient absorption spectroscopy 25, 26, our work explores a fundamentally distinct regime, the nonperturbative regime, where carrier saturation evolves within tens of femtoseconds and directly influences harmonic generation.
+This bridges the gap between weak-field saturation effects and field-driven dynamics in Dirac materials, uncovering a new class of ultrafast phenomena governed by light-driven population reshaping.
+More broadly, the ability to resolve femtosecond-scale carrier saturation through HHG suppression suggests that this approach could also serve as a sensitive, all-optical probe of light-induced phase transitions in solids as demonstrated in both experimental 27, 28, 29 and theoretical 30, 31, 32, 33 studies.
 Results
 Nonperturbative harmonic generation in HOPG
-We generate high harmonics from an HOPG sample in reflection geometry 28 using an infrared femtosecond pulse (frequency ) combined with a weak, synchronized second harmonic (SH) field (), with a polarization parallel to the fundamental pulse (see Fig.
+We generate nonperturbative harmonics from an HOPG sample in reflection geometry 34 using an infrared femtosecond pulse (frequency ) combined with a weak, synchronized second harmonic (SH) field (), with a polarization parallel to the fundamental pulse (see Fig.
 1a).
-The fundamental pump pulse has a central wavelength of 1980 nm and a duration of 60 fs (about 10 optical cycles).
+The fundamental laser pulse has a central wavelength of 1980 nm and a duration of 60 fs (about 9 optical cycles).
 The SH field is generated by focusing the fundamental field onto a -barium borate (BBO) crystal.
 Both beams are split and then recombined in an interferometric setup, achieving both spatial and temporal overlap with accurate control of the two-color (–) time delay (see Methods).
-Throughout this work, we set the SH pulse intensity at 3% of the fundamental pump intensity in both the experimental measurements and theory calculations.
+Throughout this work, we set the SH pulse intensity at 3% of the fundamental laser intensity in both experimental measurements and theory calculations.
 HOPG exhibits an AB-stacked multilayer graphene structure, resulting in a stable crystalline form whose band dispersion near the Dirac point closely resembles that of monolayer graphene.
 For comparison, ab initio calculations of the electronic band structures of both graphene and HOPG are provided in Fig.
-S11 of the Supplementary Information.
-Given the moderate pump field strength used in our experiments – limited by the damage threshold of the material – the electron dynamics are effectively confined to the vicinity of the Dirac cone (Fig.
+S12 of the Supplementary Information.
+Since our HOPG samples are undoped, the Fermi level lies close to the Dirac point (), consistent with ARPES measurements on graphite 23.
+At room temperature ( meV), any residual Fermi-level shift is small compared to the driving laser photon energy (0.63 eV).
+We therefore assume for both the experiment and the theoretical modeling.
+Given the moderate laser field strength used in our experiments – limited by the damage threshold of the material – the electron dynamics are effectively confined to the vicinity of the Dirac cone (Fig.
 1b).
-In the nonperturbative regime relevant to our work, the emission of above-bandgap harmonics is primarily governed by interband polarization processes.
+In the nonperturbative regime relevant to our work, the emission of above-bandgap harmonics is primarily governed by interband polarization processes 6.
 This behavior is further corroborated by numerical simulations based on the semiconductor Bloch equations (SBEs), which reveal a clear predominance of interband contributions in the generated harmonic spectrum (see Fig.
@@
 Here, interband tunneling produces an electron-hole pair that is driven by the laser field in momentum space until electron and hole recombine and a harmonic photon is emitted 6.
-In our experiment, non-perturbative high harmonics are generated from HOPG with spectral features observed up to the fourth order (2.5 eV photon energy) near the reported harmonic cutoff in monolayer graphene around 2.4 eV 29, 30.
+In our experiment, nonperturbative harmonics are generated from HOPG with spectral features observed up to the fifth order (3.1 eV photon energy), exceeding the previously reported harmonic cutoff in monolayer graphene of about 2.4 eV 35, 36.
 Figure 1c displays the measured harmonic spectrum obtained using the 1980 nm driving pulse with a peak intensity of 72 GW cm-2 and the admixed SH pulse.
@@
 1d).
-The third harmonic exhibits a power-law dependence of approximately , deviating from the expected perturbative scaling of and confirming the involvement of strong-field electron-hole dynamics.
-This nonlinear response arises from electron-hole acceleration and recombination processes rather than multiphoton transitions, underscoring that our measurements are firmly within the non-perturbative regime.
+At low laser intensities, H3 follows the expected perturbative scaling of , but with increasing intensity, the scaling gradually transitions to .
+In the same intensity range, the fifth harmonic (H5) emerges and exhibits a distinctly nonperturbative scaling of approximately at the highest intensity, rather than the perturbative dependence.
+These clear deviations from perturbative scaling marks the onset of NPHG 36, 35, 37.
+In this regime, the nonlinear response arises primarily from carrier acceleration and recombination rather than from conventional multiphoton transitions 35, 38.
+Moreover, the harmonic yield shows a nontrivial dependence on the laser ellipticity (Fig.
+1e), consistent with a defining signature of graphene NPHG observed at longer wavelengths in experiment and theory 35, 39, 40, 37.
+In particular, the -polarized component is strongly enhanced under elliptically polarized excitation, reaching a maximum around , reflecting strong coupling between interband and intraband currents.
+Additionally, we observe a blue shift of the harmonic spectrum under strong excitation 41, reminiscent of the spectral shifts reported in atomic HHG 42 (see Fig.
+S13 in the Supplementary Information).
 Emission time shift of harmonic generation
-Two-color spectroscopy is a powerful all-optical approach capable of retrieving the temporal dynamics of the harmonic emission all the way down to the sub-cycle scale 31.
+Two-color spectroscopy is a powerful all-optical approach capable of retrieving the temporal dynamics of the harmonic emission 43, 44.
 Here, we measure the harmonic intensity as a function of the relative time delay between the two colors (–), as shown in Fig.
@@
 2a).
-As a reference, we replace the HOPG sample with a conventional bulk gapped solid HHG sample, a ZnO crystal 5, 6, and measure the above-band gap sixth harmonic (H6) intensity as a function of .
-The measurements in ZnO are performed under the same conditions as in HOPG, except that the fundamental pump intensity is increased to 384 GW cm-2 to ensure the generation of above-band gap harmonics through the interband process.
+As a reference, we replace the HOPG sample with a conventional bulk gapped solid HHG sample, a ZnO crystal 5, 6, 45, and measure the above-band gap sixth harmonic (H6) intensity as a function of .
+The measurements in ZnO are performed under the same conditions as in HOPG, except that the fundamental laser intensity is increased to 384 GW cm-2 to ensure the generation of above-band gap harmonics through the interband process.
 The band gap of the ZnO crystal is around 3.3 eV.
@@
 S2 of the Supplementary Information.
-In each of plots shown in Fig.
+In each of the plots shown in Fig.
 2, oscillations of the harmonic signal with a frequency of 4 are visible on top of an overall Gaussian-like curve.
@@
 In contrast, the H6 signal in ZnO exhibits a symmetric delay dependence, with its maximum occurring precisely at zero delay.
-The time shift corresponding to the maximum harmonic signal is extracted by fitting the entire delay-dependent harmonic trace with a Gaussian function and identifying the peak position of the fit.
+A symmetric behavior is also found for monolayer WS2, a hexagonal gapped material (see Supplementary Information).
+For HOPG, the time shift corresponding to the maximum harmonic signal is extracted by fitting the entire delay-dependent harmonic trace with a Gaussian function and identifying the peak position of the fit.
 To gain further insight into the underlying physics, we perform theoretical simulations based on both one-dimensional (1D) and two-dimensional (2D) semiconductor Bloch equations (SBEs, see Methods in the main text and Supplementary Information for details) using the same parameters as in the experiment for both HOPG and ZnO (Fig.
@@
 S6 of the Supplementary Information).
-Under a fixed interband decoherence time of fs, both the 1D and 2D models exhibit a negative delay for the harmonic maximum (note that the negative delay in 2D simulation is sensitive to the choice of T2).
-Since the 1D SBE model reproduces the main features observed in the 2D simulations, including the negative time shift of the harmonic maximum, we adopt the simpler 1D model for the remainder of our analysis.
-The simulation results show excellent agreement with the experimental data.
+We solve the SBEs to obtain the full time-dependent microscopic populations and interband coherences induced by the laser pulse.
+In contrast to semiclassical Boltzmann equation approaches 26, we incorporate electron–electron and electron–phonon scattering via phenomenological relaxation and decoherence times , so that population redistribution and decoherence are treated at the microscopic level.
+In this framework, an out-of-equilibrium “hot” carrier distribution naturally emerges during the laser–matter interaction without explicitly describing thermalization or introducing an effective electron temperature (cf.
+41).
+Under a fixed interband decoherence time of fs, both the 1D and 2D models exhibit a negative delay for the harmonic maximum (note that the negative delay in the 2D simulation is sensitive to the choice of , see Fig.
+S11 in the Supplementary Information for details).
+Since the 1D SBE model reproduces the main features observed in the 2D simulations, including the negative time shift of the harmonic maximum, we adopt the simpler 1D model for most of the remainder of our analysis.
+The simulation results show good agreement with the experimental data.
 We accurately reproduce the time shift of the H4 signal in HOPG, yielding a similar shift fs with respect to , and correctly predict the peak of the H6 signal in ZnO at zero delay.
-Furthermore, we simulate the delay dependence of H4, H6, and H8 from ZnO, and in all cases, the harmonic maxima occur near , confirming that the choice of H6 as a reference does not affect the evaluation of the delay.
-To gain deeper insight into the pronounced time shift observed in HHG from HOPG, we analyze the time evolution of both the conduction band electron population and the interband polarization , as shown in Figs.
+Furthermore, we simulate the delay dependence of H4, H6 and H8 from ZnO, and in all cases the harmonic maxima occur near , confirming that the choice of H6 as reference does not affect the evaluation of the delay.
+In addition, supplementary calculations that coherently average the harmonic emission over the optical penetration depth in the -direction (Figs.
+S15–S16) show that depth-dependent interference does not create a negative delay in a gapped system and has only a minor effect in the gapless case, indicating that the observed temporal shift cannot be attributed to propagation or geometric averaging effects.
+To gain deeper insight into the pronounced time shift observed in NPHG from HOPG, we analyze the time evolution of both the conduction band electron population and the interband polarization , as shown in Figs.
 3a and b.
-Our calculations reveal that this shift originates from the strong interplay between intraband and interband currents in the gapless Dirac cone.
+Our calculations reveal that this time shift originates from the strong interplay between intraband and interband currents – a characteristic of NPHG in gapless Dirac materials that also underlies their nontrivial ellipticity response 46, 47, 37.
 A significant accumulation of intraband population leads to a suppression of the interband polarization.
 Notably, the conduction band population near the Dirac point rapidly approaches half-filling well before the peak of the driving field, indicating a pronounced saturation effect.
-Unlike conventional solid-state HHG processes where depletion is negligible 32, the time-dependent -space populations of conduction () and valence () bands play a crucial role in the interband transition term in the SBEs (see the Methods section, Eq.
+Unlike conventional solid-state harmonic processes where carrier saturation is negligible 48, the time-dependent -space populations of conduction () and valence () bands play a crucial role in the interband transition term in the SBEs (see the Methods section, Eq.
 (1)).
-Once both and reach 0.5, the interband transition term vanishes, effectively blocking further interband excitation – a phenomenon known as Pauli blocking.
+Once both and reach 0.5, the interband transition term vanishes, effectively blocking further interband excitation – a phenomenon known as state blocking or state-filling saturation.
 This behavior is analogous to a two-level system where population inversion is forbidden and the excitation process saturates due to balanced electron occupation across bands.
 The inability of the conduction band population to significantly exceed half-filling for most -points around the Dirac point indicates that saturation occurs early during the fundamental driving pulse.
-This saturation not only limits further excitation, but also suppresses the interband polarization, which is an essential ingredient for interband HHG.
+This saturation not only limits further excitation, but also suppresses the interband polarization, which is an essential ingredient for interband harmonic generation.
 As a result, the dominant interband term is weakened, even near the peak of the driving pulse.
@@
 It is evident that the excitation of 50% of the final total population (which remains after the end of the laser pulse) is achieved at a significantly earlier time than the peak of the driving field, underscoring the presence of strong saturation and depletion.
-In the absence of depletion, one would expect this time to coincide with the peak of the driving field due to its temporal symmetry.
-For comparison, we also plot the calculated total conduction band population dynamics in ZnO, which shows that the excitation of 50% of the final total population occurs right at the peak field, consistent with the larger band gap that prevents efficient depletion.
+In the absence of saturation, one would expect this time to coincide with the peak of the driving field due to its temporal symmetry.
+For comparison, we also plot the calculated total conduction band population dynamics in ZnO, which shows that the excitation of 50% of the final total population occurs right at the peak field, consistent with the larger band gap that prevents efficient carrier saturation.
 Furthermore, a time-frequency analysis using a Gabor transform (Fig.
-3d) shows that the maximum intensity of the H3 in HOPG driven by a single-color 2 m pulse occurs approximately 25 fs before the field maximum ().
+3d) shows that the maximum intensity of the H3 in HOPG driven by a single-color 2 m pulse occurs before the field maximum ().
+In the 1D model this advance is approximately 25 fs, while in the 2D (graphene) model it is on the order of 10 fs, in agreement with a 3D (HOPG) model (see Supplementary Information).
+We note that there is a disagreement between these simulations on the exact prediction for the shift, none of which perfectly match with the experiment.
+However, the fact that all of these levels of theory get the correct trend is strong evidence that they indeed capture the essential physics.
 A complementary Gabor transform of the harmonic emission driven by – two-color pulses at various time delays (see Fig.
@@
 This is in contrast to wide-band gap semiconductors where the non-depletion approximation is valid and interband polarization is largely unaffected by intraband population.
-In our case, the absence of a band gap in HOPG leads to substantial carrier excitation around the Dirac cone and strongly modifies the interband contribution to HHG.
-Consequently, intraband population dynamics play a critical role in shaping the overall HHG process in gapless Dirac materials such as HOPG.
-One can also expect to see a strong saturation effect in wide-bandgap semiconductor HHG as long as the pump intensity is sufficiently large.
+In our case, the absence of a band gap in HOPG leads to substantial carrier excitation around the Dirac cone and strongly modifies the interband contribution to NPHG.
+Consequently, intraband population dynamics play a critical role in shaping the overall NPHG process in gapless Dirac materials such as HOPG.
+One can also expect to see a strong saturation effect in wide-bandgap semiconductor NPHG and HHG as long as the laser intensity is sufficiently large.
 However, the low damage threshold of these materials prevents a clear observation of this effect in gapped solids.
-Nonetheless, our simulations confirm that the harmonic emission time shifts significantly earlier even in ZnO when its band gap is artificially reduced in simulations to a nearly gapless regime, as shown in Fig.
+Nonetheless, our simulations confirm that the harmonic emission time shifts significantly earlier than even in ZnO when its band gap is artificially reduced in simulations to a nearly gapless regime, as shown in Fig.
 S9 of the Supplementary Information.
 Since the emission time shift originates from dynamic population saturation, a stronger driving laser intensity is expected to induce earlier saturation during the driving process.
-To investigate this, we experimentally measure the time shift of H4 under various pump intensities, as shown in Fig.
+To investigate this, we experimentally measure the time shift of H4 under various laser intensities, as shown in Fig.
 4a.
@@
 To gain further insight into the phenomenon, we introduce an analytical approximation (Eqs.
-8–12) based on Fermi’s golden rule, assuming non-depleted excitation from the valence band to the conduction band while accounting for the influence of electron occupation in both bands on the excitation rate.
-This model also predicts a similar dependence of on the pump intensity, as indicated by the black dashed line in Fig.
+Analytical form of population dynamics–6), assuming non-depleted excitation from the valence band to the conduction band while accounting for the influence of electron occupation in both bands on the excitation rate.
+This model also predicts a similar dependence of on the laser intensity, as indicated by the black dashed line in Fig.
 4b.
@@
 5b).
-We define as the time at which half of the electrons in this momentum-space region have transitioned to the conduction band.
+We define as the time at which a quarter of the electrons in this momentum-space region have transitioned to the conduction band.
 As shown in Fig.
-5c, the absolute value of increases with pump intensity, indicating an intensity-dependent excitation process.
+5c, the absolute value of increases with laser intensity, indicating an intensity-dependent excitation process.
 Furthermore, a strong correlation is observed between (in the single-color 2 m excitation case) and (emission time shift corresponding to the maximum H4 intensity in the two-color excitation case).
 The correspondence can be captured by a second-order polynomial fitting function, , with an value of 0.99, indicating very strong correlation between these two physical observables.
+Unlike transient absorption spectroscopy, which mostly probes post-excitation relaxation dynamics, our two-color harmonic spectroscopy directly tracks the coherent, ultrafast carrier dynamics while they are driven by the strong optical field.
 This demonstrates an all-optical method for probing carrier population dynamics in Dirac materials and semi-metallic systems under strong-field excitation.
 Conclusion
-In conclusion, our joint experimental and theoretical study has established an HHG-based all-optical approach to observe and quantify ultrafast depletion effects and carrier excitation dynamics in solids, which also yields absolute timing information.
-With respect to conventional bulk materials, such as ZnO, we found that the Dirac semimetal HOPG shows strong valence band population saturation due to its gapless nature even under moderate laser intensities (), leading to half-filling in the vicinity of the Dirac point.
-We identified this effect as responsible for the observed fs time shifts and the suppression of the harmonic emission in the later part of the driving laser pulse through Pauli blocking.
-Our study demonstrates that gapless materials are subject to depletion effects and Pauli blocking, leaving behind a sizable conduction band population and rendering harmonic generation processes nonparametric in nature.
-Petahertz optoelectronics, where optical manipulation of the material is performed on attosecond time scales, relies mainly on virtual charge carrier excitations and their reversibility 33, 34, 8.
-In order to realize petahertz optoelectronics with full reversibility in gapless materials, pulse duration and peak intensity need to be limited such that Pauli blocking is avoided.
-Even if this cannot be achieved, gapless materials can nevertheless serve as an ultrafast optoelectronic switch where the presence or the absence of a pre-excitation pulse affects the efficiency of nonlinear processes through Pauli blocking on the significantly longer time scale of spontaneous emission.
-The nonparametric nature of HHG in gapless materials even at moderate intensities may also be helpful for generating upconverted light with nonclassical photon statistics due to the depletion effect, as recently predicted for gas-phase HHG 35.
+In conclusion, our joint experimental and theoretical study has established an NPHG-based all-optical approach to observe and quantify ultrafast carrier saturation effects and carrier excitation dynamics in solids, which also yields absolute timing information.
+With respect to conventional bulk materials, such as ZnO, we found that the Dirac semimetal HOPG shows strong conduction band population saturation due to its gapless nature even under moderate laser intensities (), leading to half-filling in the vicinity of the Dirac point.
+We identified this effect as responsible for the observed fs time shifts and the suppression of the harmonic emission in the later part of the driving laser pulse through state blocking.
+Our study demonstrates that gapless materials are subject to saturation effects and state blocking, leaving behind a sizable conduction band population and rendering harmonic generation processes nonparametric in nature.
+Petahertz optoelectronics, where optical manipulation of the material is performed on attosecond time scales, relies mainly on virtual charge carrier excitations and their reversibility 49, 50, 8.
+In order to realize petahertz optoelectronics with full reversibility in gapless materials, pulse duration and peak intensity need to be limited such that state blocking is avoided.
+Even if this cannot be achieved, gapless materials can nevertheless serve as an ultrafast optoelectronic switch where the presence or the absence of a pre-excitation pulse affects the efficiency of nonlinear processes through state blocking on the significantly longer time scale of spontaneous emission.
+The nonparametric nature of HHG in gapless materials even at moderate intensities may also be helpful for generating upconverted light with nonclassical photon statistics due to a saturation or depletion effect, as recently predicted for gas-phase HHG 51.
 Methods
 Experimental setup
-In the experiment, the fundamental () pulse at 1980 nm is generated from a phase-stable optical parametric chirped pulse amplification (OPCPA) laser system operating at a repetition rate of 1 MHz (Class 5 Photonics White Dwarf 36).
+In the experiment, the fundamental () pulse at 1980 nm is generated from a phase-stable optical parametric chirped pulse amplification (OPCPA) laser system operating at a repetition rate of 1 MHz (Class 5 Photonics White Dwarf 52).
 The pulse duration of the field measured by frequency-resolved optical gating (FROG) yields a full-width-at-half-maximum (FWHM) duration of approximately 60 fs ( optical cycles).
@@
 The relative delay between the two pulses was controlled via a piezo delay stage placed in the beam path.
-The two beams were then recombined by reflecting them off two mirrors back into the same dichroic beam splitter, with a slight lateral offset angle relative to the incoming paths, and subsequently directed toward the sample for HHG.
+The two beams were then recombined by reflecting them off two mirrors back into the same dichroic beam splitter, with a slight lateral offset angle relative to the incoming paths, and subsequently directed toward the sample for harmonic generation.
 A parabolic mirror with mm focuses the light onto the HOPG sample (TipsNano) at an angle of incidence of 14 degrees.
@@
 Semiconductor Bloch equations calculations
-To describe the interaction of the two-color field and HOPG in the Dirac cone, we use the following one-dimensional semiconductor Bloch equations37, 38 (SBEs) centered around the Dirac cone:
+To describe the interaction of the two-color field and HOPG in the Dirac cone, we use the following one-dimensional semiconductor Bloch equations53, 54 (SBEs) centered around the Dirac cone:
 Here, represents a dimensionless polarization depending on time and momentum , induced by the time-dependent electric field .
@@
 This approximation should be valid in our regime where the laser-driven electrons only move near the Dirac cone.
-The transition dipole moments (TDMs) are calculated as 39 (see Fig.
+The transition dipole moments (TDMs) are calculated as 46 (see Fig.
 S3 in the Supplementary Information).
 We set the phenomenological interband polarization decoherence time to one optical cycle of the 2 m driving laser field fs, and the relaxation time for intraband current to .
-This consideration is valid since the relaxation is generally long and ranges from few hundred femtosecond to picosecond in graphene and is less sensitive to the results 40, 41.
+This consideration is valid since the relaxation is generally long and ranges from few hundred femtosecond to picosecond in graphene and is less sensitive to the results 47, 39.
 Band structure and TDM for ZnO crystal along the –M direction are shown in Fig.
@@
 The yield of the harmonics for the two-color case are calculated with two-color delay steps of 0.33 fs.
-The HHG spectrum is computed as 37
+The HHG spectrum is computed as 53
 where
 and
-denote the interband and intraband contributions to the HHG yield, respectively. represents the group velocity of the electrons and holes in the corresponding bands.
+denote the interband and intraband contributions to the harmonic yield, respectively. represents the group velocity of the electrons and holes in the corresponding bands.
 Analytical form of population dynamics
 For a theoretical analysis of saturation and electron population dynamics, we begin by considering the non-depleted excitation rate of electrons from the valence band to the conduction band.
-According to Fermi’s golden rule, this excitation rate at a given crystal momentum and time can be expressed as
-The total excitation rate integrated over the Brillouin zone is then given by
+The excitation rate at a given crystal momentum and time can be expressed as
+w_0(k, t) ∼2πℏ — ⟨ψ_c — E(t) ⋅^d — ψ_v ⟩—^2 ≈2π— E(t) ⋅d_cv(k+A(t)) —^2.
+The total excitation rate integrated over the Brillouin zone is then given by {align} w(t) = ∫w_0(k, t) dk = 2πE^2(t) ∫—d_cv(k+A(t))—^2 dk.
 Let denote the total conduction band electron population at time . The evolution of this population in discrete time steps can be described by
+P(t_n+1) = P(t_n) + (1 - P(t_n)) w(t_n)dt - P(t_n) w(t_n)dt,
+P(tn+1) - P(tn)1 - 2P(tn) = w(t_n)dt.
 Transitioning to the continuous-time limit, we obtain the differential form
+{align}
+dP(t)/(1 - 2P(t)) = w(t)dt,
+dln(1 - 2P(t)) = -2w(t)dt.
 Integrating yields the electron population as a function of time,
 The actual (depleted) instantaneous excitation rate is given by
-For interband-driven high-harmonic generation (HHG), the emitted harmonic intensity is proportional to the instantaneous excitation rate, .
+For interband-driven harmonic generation, the emitted harmonic intensity is proportional to the instantaneous excitation rate, .
 Data availability
@@
 G.
+and E.
+U.
 and O.
@@
 C.
-and C.
-G.
 wrote the initial manuscript.
@@
 Corresponding authors
-Correspondence to Michael Krüger or Marcelo F. Ciappina.
+Correspondence to Michael Krüger or Marcelo F. Ciappina or Ofer Neufeld.
 Acknowledgements.
@@
 Yan and Y.
-Wu for insightful discussions and U.
+Wu for insightful discussions, U.
 Leonhardt and L.
 M.
-Procopio for providing specialized equipment.
+Procopio for providing specialized equipment, and X.
+Duan for providing the monolayer WS2 sample.
 Z.
@@
 F.
-C and M.
+C.
+and M.
 K.
@@
 N., D.
-K., M.
-F.
-C and M.
+K., E.
+U., O.
+N.
+and M.
 K.
-thank the Helen Diller Quantum Center at the Technion for partial financial support.
+thank the Helen Diller Quantum Center and the Russell Berrie Nanotechnology Institute at the Technion for partial financial support.
 M.
@@
 acknowledges support by the National Key Research and Development Program of China (Grant No.
-2023YFA1407100), Guangdong Province Science and Technology Major Project (Future functional materials under extreme conditions - 2021B0301030005) and the Guangdong Natural Science Foundation (General Program project No.
-2023A1515010871).Ethics declarations
+2023YFA1407100), Guangdong Province Science and Technology Major Project (Future functional materials under extreme conditions - 2021B0301030005), the Guangdong Natural Science Foundation (General Program project No.
+2023A1515010871) and the National Natural Science Foundation of China (Grant No.
+12574092).Ethics declarations
 The authors declare no competing interests.